[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Fri Aug 11, 2017

TAFB Shift Log nhc.tsbadmin at noaa.gov
Sat Aug 12 02:04:51 UTC 2017


TAFB Shift Log for Fri Aug 11 2017 22:04:45 GMT-0400 (EDT)

Shift: Swing Shift

Pacific Forecaster: J Lewitsky

Atlantic Forecaster: D Mundell

Surface Analyst: M Formosa

Technical Issues Summary:

GFE/NCP issues: JL and Pablo tested the Pacific NWPS throughout the shift.  
Still troubleshooting.

N-AWIPS issues:

Other issues:

Miscellaneous issues: Received a ADMNCF message in AWIPS about NCF learning  
that the Gateway is not passing a variety of products via SBN uplink (saw a  
momentary drop in obs at NHC). Also received a message that all systems at  
Boulder have been moved to College park due to an outage. SDM called at  
430pm to tell us about Critical Weather day from 00z Sat - 12z Mon for  
heavy rainfall over the central US (TX/OK). Another MIAADMNFD message was  
received at 0125z saying that no upper air data for 00z was ingested. MRF  
notified the WFO MFL who did not receive the message.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170812/35bb7de0/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list